OpenStudy (anonymous):

Combine like terms. (7y^2 - 3y) + 8y^2 - 8y)

OpenStudy (anonymous):

The like terms are: 7y^2 and 8y^2 -3y and -8y. Add them. 7y^2 + 8y^2 = 15y^2 -3x + (-8y) = -3y - 8y = -11y So the answer would be 15y^2 - 11y .

OpenStudy (anonymous):

you're so smart! @jram2326 thanks:)

OpenStudy (anonymous):

You're welcome @XxNina26xX :)

OpenStudy (anonymous):

@jram2326 give me a medal..lol :)

OpenStudy (anonymous):

thanks :p

OpenStudy (anonymous):

There :) @XxNina26xX